# STP: Self-play LLM Theorem Provers with Iterative Conjecturing and Proving

### v2

@@ Abstract
A fundamental challenge in formal theorem proving by LLMs is the lack of high-quality training data. Although reinforcement learning or expert iteration partially mitigates this issue by alternating between LLM generating proofs and finetuning them on correctly generated ones, performance quickly plateaus due to the scarcity of correct proofs (sparse rewards). To keep improving the models with limited data, we draw inspiration from mathematicians, who continuously develop new results, partly by proposing novel conjectures or exercises (which are often variants of known results) and attempting to solve them. We design the Self-play Theorem Prover (STP) that simultaneously takes on two roles, conjecturer and prover, each providing training signals to the other. The conjecturer is trained iteratively on previously generated conjectures that are barely provable by the current prover, which incentivizes it to generate increasingly challenging conjectures over time. The prover attempts to prove the conjectures with standard expert iteration. We evaluate STP with both Lean and Isabelle formal versifiers. With 19.8 billion tokens generated during the training in Lean, STP proves 26.3% of the statements in the LeanWorkbook dataset, doubling the previous best result of 13.2% achieved through expert iteration. The final model achieves state-of-the-art performance among whole-proof generation methods on miniF2F-test (61.1%, pass@3200), Proofnet-test (23.1%, pass@3200) and PutnamBench (8/644, pass@64).
@@ 1 Introduction
The reasoning capability of large language models (LLMs) is critical for various applications, including coding assistants, question-answering, and agents (Plaat et al., 2024; Shinn et al., 2023; Yao et al., 2022; Shao et al., 2024; Li et al., 2023; Nijkamp et al., 2022). It is also a key criterion for achieving artificial general intelligence (AGI). Automated theorem proving with formal languages by LLMs stands at the forefront of reasoning research, partly because it allows objective and reliable evaluation through classical verifiers such as Lean (Moura and Ullrich, 2021) and Isabelle (Nipkow et al., 2002). Moreover, it arguably encapsulates the essence of advanced reasoning tasks while abstracting away the ambiguity of natural language, enabling meaningful studies on a relatively smaller scale.
@@
Reinforcement learning (RL) on datasets without solutions (e.g., datasets with theorem statements or reasoning questions and answers) is a prominent approach for improving the reasoning capability, as seen in the recent development of OpenAI o1 Jaech et al. (2024), DeepSeek-Prover (Xin et al., 2024a) and DeepSeek R1 (Guo et al., 2025). Often referred to as expert iteration Anthony et al. (2017), it partially mitigates the data scarcity issue by alternating between LLMs generating proofs and finetuning them on correctly generated ones (Kaliszyk et al., 2018; Wu et al., 2021; AlphaProof, 2024; Xin et al., 2024b; Ying et al., 2024).
However, as Wu et al. (2024) pointed out, RL or expert iteration often saturates at a low pass rate because the number of samples required to generate a correct proof for an unproven theorem grows exponentially. As a result, a massive amount of computation is wasted on generating incorrect proofs that provide no training signal to the model. For instance, in the proof sampling process of Wu et al. (2024), 98.5% of the compute yields no successful proofs, despite the pass rate being only 13.2% on LeanWorkbook. In other words, after a few rounds of expert iteration, re-training the model becomes much less effective due to the limited number of new successful proofs.
@@
Figure 3: Comparison of pass rates on miniF2F-test (y-axis) with different numbers of inference-time samples (x-axis). The model trained with STP consistently outperforms the DeepSeek-Prover-V1.5 series.
We empirically evaluate our method with both Lean (Moura and Ullrich, 2021) and Isabelle (Nipkow et al., 2002). For the Lean experiments, we aim for the best performance and therefore choose DeepSeek-Prover-V1.5-SFT (Xin et al., 2024b) as the base model for STP. As shown in Fig. 3, after a self-play training of roughly 120M generated proofs and 2M generated conjectures, we successfully prove 26.3% of the statements in the training dataset LeanWorkbook (Ying et al., 2024), doubling the previous best result of 13.2% (Wu et al., 2024) achieved by expert iteration. In Fig. 3, we compare the inference-time performance of existing models and the final model trained with STP by taking multiple independent samples on a common benchmark, miniF2F-test (Zheng et al., 2021). Our model significantly outperforms the DeepSeek-Prover-V1.5 models across various sampling budgets. We also achieve state-of-the-art performance among whole-proof generation methods on miniF2F-test (61.1%, pass@3200), ProofNet-test (23.1%, pass@3200) (Azerbayev et al., 2023a) and PutnamBench (8/644, pass@64) (Tsoukalas et al., 2024), where pass@k represents the percentage of statements proved with $k$ independently sampled proofs per statement.
In the Isabelle experiments, we study the scalability of STP by starting from a generic math-focused model Llemma-7b (Azerbayev et al., 2023b) and run STP for many more iterations (300M generated proofs in total). We compare the scaling of STP with expert iteration and parallel sampling, by taking several model checkpoints during the STP training run and then switching to the baseline methods. The results clearly demonstrate that STP achieves a better scaling behavior starting from various checkpoints with different capability (see Fig. 4 (Left) in Section 4.3). Ablation study also demonstrates that the main performance gain stems from the dense training signals given by the conjectures. Expert iteration wasted its compute on generating unsuccessful proofs to challenging theorems in the training dataset—at a checkpoint where the pass rate is around 11.4% on LeanWorkbook, only 131 out of 2.5M generated proofs of the unproved statements are correct, resulting in very limited training signals. In contrast, at least 47% of the generated conjectures in STP training are successfully proved because the conjecturer is trained to generate more approachable statements thanks to the design of its reward (see Fig. 4 (Right)).
@@ 2 Additional Related Works
The closest related work to this paper is the concurrent work of Poesia et al. (2024) which also designs a self-play training that iterates between conjecturing and theorem proving. The key difference between this paper and Poesia et al. (2024) is that we start with a pre-trained model and work on practical formal languages like Lean and Isabelle with an infinite space of possible proof steps (which are actions in the RL algorithm), whereas Poesia et al. (2024) operates in a simplified and constrained setting with a finite action space and trains from scratch. As a result, Poesia et al. (2024) rely on constrained decoding to force the validity of generated conjectures, while we solely rely on the LLM itself to generate valid conjectures. Technically, since our training process is much longer (more than 50 iterations) than Poesia et al. (2024) (5 iterations), we must carefully design the conjecturing reward to maintain the diversity and relevance of the generated conjectures (see Section 3.2).
The idea of generating new tasks by the model is also explored in other domains such as alignment (Ye et al., 2024), programming puzzles (Haluptzok et al., 2022; Teodorescu et al., 2023; Pourcel et al., 2024b), video games (Zhang et al., 2023; Pourcel et al., 2024a), and classic RL environments (Parker-Holder et al., 2022; Colas et al., 2022).
More generally, self-play training has demonstrated its potential to achieve super-human performance on two-player games in a fixed environment like Go (Silver et al., 2016).
@@ 3.2 Self-play training
We use weighted cross entropy loss computed on the conjectures or proofs (but not the inputs of the model). For the proof dataset, we weight the examples reciprocally to the number of verified proofs to the corresponding statement/conjecture. We also use a length penalization of the form $\gamma^{L}$ to reward simpler proofs, where $\gamma<1$ is the discount factor and $L$ is the length of the proof.
@@ 4.1 Implementation details
We use the de-duplicated LeanWorkbook (Ying et al., 2024), which contains around 89K Lean4 statements, as the given dataset without proofs (see Appendix A.4 for details). For the Isabelle experiments, we translate the Lean4 statements to Isabelle using the DeepSeek V2.5 with few-shot prompting.
@@
For inference, we cap the number of generated tokens to 1024, and set the sampling temperature to 0.7 for Llemma-7b and 1.0 for DeepSeek-Prover, following Dong et al. (2024); Xin et al. (2024b), respectively. For training, we use batch size 2048 and Adam (Kingma and Ba, 2014) with a constant learning rate of 5e-5 in STP, and 1e-4 in SFT and final re-training. The discount factor is $\gamma=\exp(-0.001).$
@@ 4.2 Results with Lean
For the Lean experiments, we choose DeepSeek-Prover-V1.5-SFT as our base model, which was trained on proofs collected by expert iteration on a combination of public, such as LeanWorkbook, miniF2F-valid (Zheng et al., 2021), and ProofNet-valid (Azerbayev et al., 2023a), and other proprietary datasets. We ran 24 iterations of STP and generated 2M conjectures, 120M proofs, and 19.8B tokens in total. We use the cumulative pass rate, defined by the fraction of statements proved during the entire training, as the main metric for training progress. Fig. 3 plots the cumulative pass rate of STP and two major baselines, expert iteration, and parallel sampling, on the training dataset LeanWorkbook (Ying et al., 2024). Expert iteration alternates between generating proofs to the statements in the given dataset and finetuning the model on correct proofs. (See discussions and comparison about variants of expert iteration in Appendix A.6.) Parallel sampling simply generates proofs with the given model. Fig. 3 shows that STP achieves significantly better scaling than expert iteration, which simulates the performance of DeepSeek’s model as if it were trained for more iterations.
To achieve the best performance on common benchmarks, we further train the model using statements from LeanWorkbook, miniF2F-valid, and ProofNet-valid for another 8 iterations. Table 1 compares our models with prior works on miniF2F-test and ProofNet-test, which contain formal statements of high-school level and college level math questions, respectively. Among the whole-proof generation methods, STP significantly outperforms DeepSeek-Prover-V1.5-RL (which is continuously trained with RL on top of their SFT model) and achieves SoTA performance across various inference-time sample budgets.
@@
3200 $28.2$K 61.1% 23.1%
STP    (w/ miniF2F-valid, ProofNet-valid)  [^4] Our base model, DeepSeek-Prover-V1.5-SFT, is trained on miniF2F-valid and ProofNet-valid, although we only include these statements in the last 8 iterations of this experiment. 128 1.1K 57.7% $\pm$ 0.6% 18.5% $\pm$ 0.8%
3200 28.0K 61.7% $\pm$ 0.6% 23.1% $\pm$ 0.5%
25,600 224K 63.5% 25.8%
@@
As shown in Table 3, on PutnamBench (Tsoukalas et al., 2024) which consists of undergraduate-level mathematics competition questions, STP solves 8 out of 644 problems with 64 samples per problem, and STP (w/ miniF2F valid, ProofNet valid) solves 9 problems with 3200 samples per problem, outperforming the best result of 6 problems in prior works achieved by Wu et al. (2024).
@@ 4.4 Ablation study
One may hypothesis that the self-play algorithm and generated conjectures only help improve the pass rate on LeanWorkBook. It turns out that in the final re-training stage, it is still beneficial to re-train with the generated conjectures in addition to the successfully proved statements in LeanWorkBook even for performance on miniF2F-test and ProofNet-test—it leads to about 1% performance gain (for pass@128) than re-training only on the latter (See Appendx B.1).
@@ 4.5 Examples of generated conjectures
theorem lean_workbook_54038 (x : ℕ) (n : ℕ) (hn : 1 < n) : (x^n - 1) %
@@ A.5 Re-weighting the conjecturing dataset
In the actual implementation, we additionally requires that the weighting $P$ for every conjecture $x$ cannot be too big because otherwise it might cause instability of the LLM training with weighted cross entropy loss. Therefore, our practical implementation is shown in Alg. 4.
@@
5:     Compute $x^{\star}=\operatorname*{argmin}_{x\in{\mathcal{X}}}d(x,y)M(x).$
6:     $P(x^{\star})\leftarrow P(x^{\star})+1/m.$
7:     if $P(x^{\star})*n>3$ then
8:         $M(x^{\star})\leftarrow 0.$
9:Return: the optimal re-weighting is $[P(x_{1})*n,P(x_{2})*n,\cdots,P(x_{n})*n].$
@@
Figure 5: Left: Comparison of pass rates between STP, two implementations of expert iteration, and parallel sampling methods on LeanWorkbook. Right: Comparison of pass rates between STP and baseline methods on LeanWorkbook, starting with the SFT model. The red crosses shows the points where we refresh the self-play training as described in Section 4.1.

### v4

@@ Abstract
A fundamental challenge in formal theorem proving by LLMs is the lack of high-quality training data. Although reinforcement learning or expert iteration partially mitigates this issue by alternating between LLM generating proofs and finetuning them on correctly generated ones, performance quickly plateaus due to the scarcity of correct proofs (sparse rewards). To keep improving the models with limited data, we draw inspiration from mathematicians, who continuously develop new results, partly by proposing novel conjectures or exercises (which are often variants of known results) and attempting to solve them. We design the Self-play Theorem Prover (STP) that simultaneously takes on two roles, conjecturer and prover, each providing training signals to the other. The conjecturer is trained iteratively on previously generated conjectures that are barely provable by the current prover, which incentivizes it to generate increasingly challenging conjectures over time. The prover attempts to prove the conjectures with standard expert iteration. We evaluate STP with both Lean and Isabelle formal versifiers. With 51.3 billion tokens generated during the training in Lean, STP proves 28.5% of the statements in the LeanWorkbook dataset, doubling the previous best result of 13.2% achieved through expert iteration. The final model achieves state-of-the-art performance among whole-proof generation methods on miniF2F-test (65.0%, pass@3200), ProofNet-test (23.9%, pass@3200) and PutnamBench (8/644, pass@3200). We release our code, model, and dataset in this url: https://github.com/kfdong/STP.
@@ 1 Introduction
The reasoning capability of large language models (LLMs) is critical for various applications, including coding assistants, question-answering, and agents (Plaat et al., 2024; Shinn et al., 2023; Yao et al., 2022; Shao et al., 2024; Li et al., 2023; Nijkamp et al., 2022). It is also a key criterion for achieving artificial general intelligence (AGI). Automated theorem proving with formal languages by LLMs stands at the forefront of reasoning research (Yang et al., 2024a), partly because it allows objective and reliable evaluation through classical verifiers such as Lean (Moura and Ullrich, 2021) and Isabelle (Nipkow et al., 2002). Moreover, it arguably encapsulates the essence of advanced reasoning tasks while abstracting away the ambiguity of natural language, enabling meaningful studies on a relatively smaller scale.
@@
Reinforcement learning (RL) on datasets without solutions (e.g., datasets with theorem statements or reasoning questions and answers) is a prominent approach for improving the reasoning capability, as seen in the recent development of OpenAI o1 Jaech et al. (2024), DeepSeek-Prover (Xin et al., 2024a) and DeepSeek R1 (Guo et al., 2025). Often referred to as expert iteration (Anthony et al., 2017), it partially mitigates the data scarcity issue by alternating between LLMs generating proofs and finetuning them on correctly generated ones (Kaliszyk et al., 2018; Wu et al., 2021; AlphaProof, 2024; Xin et al., 2024b; Ying et al., 2024).
However, as Wu et al. (2024) pointed out, RL or expert iteration often saturates at a low pass rate because the number of samples required to generate a correct proof for an unproven theorem grows exponentially. As a result, a massive amount of computation is wasted on generating incorrect proofs that provide no training signal to the model. For instance, in the proof sampling process of Wu et al. (2024), 98.5% of the compute yields no successful proofs, despite the pass rate being only 13.2% on the training dataset, LeanWorkbook (Ying et al., 2024). In other words, after a few rounds of expert iteration, re-training the model becomes much less effective due to the limited number of new successful proofs.
@@
Refer to caption Figure 3: Comparison of pass rates on miniF2F-test (y-axis) with different numbers of inference-time samples (x-axis). The model trained with STP consistently outperforms the DeepSeek-Prover-V1.5 series.
We empirically evaluate our method with both Lean (Moura and Ullrich, 2021) and Isabelle (Nipkow et al., 2002). For the Lean experiments, we aim for the best performance and therefore choose DeepSeek-Prover-V1.5-SFT (Xin et al., 2024b) as the base model for STP. As shown in Fig. 3, after a self-play training of roughly 241M generated proofs and 3.6M generated conjectures, we successfully prove 28.5% of the statements in the training dataset LeanWorkbook (Ying et al., 2024), doubling the previous best result of 13.2% (Wu et al., 2024) achieved by expert iteration. In Fig. 3, we compare the inference-time performance of existing models and the final model trained with STP by taking multiple independent samples on a common benchmark, miniF2F-test (Zheng et al., 2021). Our model significantly outperforms the DeepSeek-Prover-V1.5 models across various sampling budgets. We also achieve state-of-the-art performance among whole-proof generation methods on miniF2F-test (65.0%, pass@3200), ProofNet-test (23.9%, pass@3200) (Azerbayev et al., 2023a) and PutnamBench (8/644, pass@3200) (Tsoukalas et al., 2024), where pass@k represents the percentage of statements proved with $k$ independently sampled proofs per statement.
In the Isabelle experiments, we study the scalability of STP by starting from a generic math-focused model Llemma-7b (Azerbayev et al., 2023b) and run STP for more iterations (300M generated proofs in total). We compare the scaling of STP with expert iteration and parallel sampling, by taking several model checkpoints during the STP training run and then switching to the baseline methods. The results clearly demonstrate that STP achieves a better scaling behavior starting from various checkpoints with different capability (see Fig. 4 (Left) in Section 4.3). Ablation study also demonstrates that the main performance gain stems from the dense training signals given by the conjectures. Expert iteration wasted its compute on generating unsuccessful proofs to challenging theorems in the training dataset—at a checkpoint where the pass rate is around 11.4% on LeanWorkbook, only 131 out of 2.5M generated proofs of the unproved statements are correct, resulting in very limited training signals. In contrast, at least 47% of the generated conjectures in STP training are successfully proved because the conjecturer is trained to generate more approachable statements thanks to the design of its reward (see Fig. 4 (Right)).
@@ 2 Additional Related Works
The closest related work to this paper is Poesia et al. (2024) which also designs a self-play training that iterates between conjecturing and theorem proving. The key difference between this paper and Poesia et al. (2024) is that we start with a pre-trained model and work on practical formal languages like Lean and Isabelle with an infinite space of possible proof steps (which are actions in the RL algorithm), whereas Poesia et al. (2024) operates in a simplified and constrained setting with a finite action space and trains from scratch. As a result, Poesia et al. (2024) rely on constrained decoding to force the validity of generated conjectures, while we solely rely on the LLM itself to generate valid conjectures. Technically, since our training process is much longer (more than 50 iterations) than Poesia et al. (2024) (5 iterations), we must carefully design the conjecturing reward to maintain the diversity and relevance of the generated conjectures (see Section 3.2).
The idea of generating new tasks by the model is also explored in other domains such as alignment (Ye et al., 2024), programming puzzles (Haluptzok et al., 2022; Teodorescu et al., 2023; Pourcel et al., 2024b), video games (Zhang et al., 2023; Pourcel et al., 2024a), and classic RL environments (Parker-Holder et al., 2022; Colas et al., 2022). More generally, self-play training has demonstrated its potential to achieve super-human performance on two-player games in a fixed environment like Go (Silver et al., 2016).
@@ 3.2 Self-play training
We use weighted cross entropy loss computed on the conjectures or proofs (but not the inputs of the model). For the proof dataset, we weight the examples reciprocally to the number of verified proofs to the corresponding statement/conjecture. We also use a length penalization of the form $\gamma^{L}$ to reward simpler proofs, where $\gamma<1$ is the discount factor and $L$ is the length of the proof. For the experiments with Lean, we additionally reward proofs that has faster verification time by a penalization of the form $\beta^{T}$, where $T$ is the execution time of the Lean verifier. [^3] In our preliminary experiments, we found that without the penalization on verification time, the Lean verifier takes 2x more wall-clock time on CPU than sampling proofs on TPU for our cluster setup, which becomes a bottleneck for STP training.
@@ 4.1 Implementation details
Our primary source of statements without proofs is the de-duplicated LeanWorkbook (Ying et al., 2024), which contains around 89K Lean4 statements (see Appendix A.4 for details). For the Isabelle experiments, we translate the Lean4 statements to Isabelle using the DeepSeek V2.5 with few-shot prompting. For the Lean experiments, we combine LeanWorkbook, miniF2F-valid, and ProofNet-valid as the training dataset for STP.
@@
For inference, we cap the number of generated tokens to 1024, and set the sampling temperature to 0.7 for Llemma-7b and 1.0 for DeepSeek-Prover, following Dong et al. (2024); Xin et al. (2024b), respectively. For training, we use batch size 2048 and Adam (Kingma and Ba, 2014) with a constant learning rate of 5e-5 in STP, and 1e-4 in SFT and final re-training. The discount factors are $\gamma=\exp(-0.001)$ and $\beta=\exp(-0.01)$
@@ 4.2 Results with Lean
For the Lean experiments, we choose DeepSeek-Prover-V1.5-SFT as our base model, which is trained on proofs collected by expert iteration on a combination of public, such as LeanWorkbook, miniF2F-valid (Zheng et al., 2021), and ProofNet-valid (Azerbayev et al., 2023a), and proprietary datasets. We run 48 iterations of STP and generated 3.6M conjectures, 241M proofs, and 51.3B tokens in total. We use the cumulative pass rate, defined by the fraction of statements proved during the entire training, as the main metric for training progress. Fig. 3 plots the cumulative pass rate of STP and two major baselines, expert iteration, and parallel sampling, on the training dataset LeanWorkbook (Ying et al., 2024). Expert iteration alternates between generating proofs to the statements in the given dataset and finetuning the model on correct proofs. (See discussions and comparison about variants of expert iteration in Appendix A.6.) Parallel sampling simply generates proofs with the given model. Fig. 3 shows that STP achieves significantly better scaling than expert iteration, which simulates the performance of DeepSeek’s model as if it were trained for more iterations.
Since the formal statements in our training dataset, LeanWorkbook, are translated from natural language statements, they are not always provable. In Appendix B.3, we randomly select 20 unproved statements from LeanWorkbook and manually assess whether (a) the formal statement is an accurate translation of the natural language statement, and (b) the formal statement itself is correct and provable. We find that 16 out of the 20 statements are translated correctly, but only 7 statements are provable and the remaining 13 statements are unprovable (e.g., due to missing assumptions in the corresponding natural language statement), suggesting that the best possible pass rate on LeanWorkbook, with a 95% confidence interval, is between 38.7% and 68.5%.
In Table 1, we compare the final re-trained model of STP with prior works on two common benchmarks, miniF2F-test and ProofNet-test, which contain formal statements of high-school level and college level math questions, respectively. Among the whole-proof generation methods, STP significantly outperforms DeepSeek-Prover-V1.5-RL (which is continuously trained with RL on top of their SFT model) and achieves SoTA performance across various inference-time sample budgets. We also report the performance of the model trained only on LeanWorkbook for 24 iterations, excluding miniF2F-valid and proofnet-valid, demonstrating that the model trained with STP also generalizes to out-of-domain theorems. [^5] Our base model, DeepSeek-Prover-V1.5-SFT, is trained on miniF2F-valid and ProofNet-valid, though we only run STP on LeanWorkbook in this experiment. The penalization on verification time is also not included in this experiment.
@@
      (w/o miniF2F-valid, ProofNet-valid) 3200 $28$K 61.1% 23.1%
STP 128 1.3K 61.2% $\pm$ 0.6% 19.5% $\pm$ 0.7%
3200 32K 65.0% $\pm$ 0.5% 23.9% $\pm$ 0.6%
25,600 254K 67.6% 26.9%
@@
As shown in Table 3, on PutnamBench (Tsoukalas et al., 2024) which consists of undergraduate-level mathematics competition questions, STP solves 7 out of 644 problems with 128 samples per problem, and 8 problems with 3200 samples per problem, outperforming the best result of 6 problems in prior works achieved by Wu et al. (2024).
@@ 4.4 Ablation study
One may hypothesis that the self-play algorithm and generated conjectures only help improve the pass rate on LeanWorkbook. It turns out that in the final re-training stage, it is still beneficial to re-train with the generated conjectures in addition to the successfully proved statements in LeanWorkbook even for performance on miniF2F-test and ProofNet-test—it leads to about 2-3% performance gain (for pass@128) than re-training only on the latter (See Appendx B.1).
@@ 4.5 Examples of generated conjectures
theorem lean_workbook_54038 (x : ℕ) (n : ℕ) (hn : 1 < n) : (x^n - 1) % (x - 1) 1
@@ A.5 Re-weighting the conjecturing dataset
Our practical implementation is shown in Alg. 4. In this implementation, we additionally requires that the weighting $P$ for every conjecture $x$ cannot be too big because otherwise it might cause instability of the LLM training with weighted cross entropy loss. We also allow unproved statements in $Q$ to have different matching weights — an important statement can be matched to more than one conjecture (see Line 5-6 of Alg. 4). In both the Isabelle and Lean experiments, the statements from LeanWorkbook have matching weight 1. The statements from miniF2F-valid and ProofNet-valid have matching weight 1 for the first 24 iterations in the Lean experiment, and 128 afterward.
@@
5:     Let $k$ be the matching weight of $y$.
6:     Let $x^{1},\cdots,x^{k}$ be the $k$ conjectures with smallest value of $d(\cdot,y)M(\cdot).$
7:     for $i=1,\cdots,k$ do
8:         $P(x^{i})\leftarrow P(x^{i})+1/m.$
9:         if $P(x^{i})*n>3$ then
10:              $M(x^{i})\leftarrow 0.$
11:Return: the optimal re-weighting is $[P(x_{1})*n,P(x_{2})*n,\cdots,P(x_{n})*n].$
@@
Figure 5: Left: Comparison of pass rates between STP, two implementations of expert iteration, and parallel sampling methods on LeanWorkbook. Right: Comparison of pass rates between STP and baseline methods on LeanWorkbook (Isabelle translation). The red crosses shows the points where we refresh the self-play training as described in Section 4.1.